OpenStudy (anonymous):

Matt sells burgers and sandwiches. The daily cost of making burgers is $520 more than the difference between the square of the number of burgers sold and 30 times the number of burgers sold. The daily cost of making sandwiches is modeled by the following equation: C(x) = 2x2 - 40x + 300 C(x) is the cost in dollars of selling x sandwiches. Which statement best compares the minimum daily cost of making burgers and sandwiches?

OpenStudy (anonymous):

It is greater for sandwiches than burgers because the approximate minimum cost is $250 for burgers and $292 for sandwiches. It is greater for sandwiches than burgers because the approximate minimum cost is $100 for burgers and $295 for sandwiches. It is greater for burgers than sandwiches because the approximate minimum cost is $295 for burgers and $100 for sandwiches. It is greater for burgers than sandwiches because the approximate minimum cost is $292 for burgers and $250 for sandwiches.

OpenStudy (anonymous):

I don't know how to do this at all. Can someone please help with this problem ?

OpenStudy (anonymous):

Okay give me a minute, @whpalmer4 I don't know how well I can explain without giving the answer, could u help out w/ explanation :D

OpenStudy (anonymous):

Thank you

OpenStudy (whpalmer4):

We're given the formula for the cost of selling \(x\) sandwiches: \[C(x) = 2x^2-40x+300\] We're told the daily cost of making burgers is $520 more than the difference between the square of the number of burgers sold and 30 times the number of burgers sold. Let's translate the words into an expression: "$520 more" -> "+520" "square of the number of burgers sold": x^2 "30 times the number of burgers": 30x difference: - so daily cost of making burgers is \[B(x) = x^2 -30x + 520\]

OpenStudy (whpalmer4):

So now we have two equations to compare. We want to know the minimum cost for each one. Given that these formulas depend on the number of burgers or sandwiches made/sold, won't that minimum be when the number made/sold is 0? Evaluate each of the two equations with \(x=0\) to find the minimum costs. Note that the numbers in the answers may be slightly different (just to make it a bit harder to guess the answers without doing the problem, I suspect).

OpenStudy (anonymous):

So, the answer is: It is greater for burgers than sandwiches because the approximate minimum cost is $292 for burgers and $250 for sandwiches because it is about 1/2 the total when I make x=0 ?

OpenStudy (anonymous):

I'm very confused.

OpenStudy (whpalmer4):

The cost of making 0 sandwiches is\[C(0) = 2(0)^2-40(0)+300 = \] The cost of making 0 burgers is \[B(0) = (0)^2-30(0)+520 = \]

OpenStudy (whpalmer4):

Are you sure you've copied the problem and answer choices accurately?

OpenStudy (anonymous):

so it's 300 and 520 but that is not one of my options

OpenStudy (whpalmer4):

Right, I'm starting to understand your confusion here :-)

OpenStudy (anonymous):

Yes, I just checked it again.

OpenStudy (whpalmer4):

The problem is also sloppily written, in my opinion, because it can't make up its mind about whether the formulas describe the cost of making or the cost of selling...

OpenStudy (whpalmer4):

I'd say the last answer choice is the closest one to being correct, although one of the numbers is wrong.

OpenStudy (whpalmer4):

Ask your teacher about it if you get a chance, but I wouldn't lose any sleep over it. Problem authors make mistakes, too.

OpenStudy (anonymous):

Thank you! :)